Two particles carrying equal charges move parallel to each other with speed 150 km/s. If F(1) and F(2) are magnetic and electric forces between two charged particles then the ratio of F(1)/ F(2) is
Ans 25×10^-8
1
Expert's answer
2020-01-24T08:22:59-0500

"q_1=q_2"

"v_1=v_2=v=150\times 10^3 \; m/s"

"c^2=8.98 \times 10^{16} \; m^2/s^2"

"F_1" magnetic force

"F_2" electric force


"\frac{F_1}{F_2} -?"


Solution:

Suppose that "q_1>0, q_2>0" ( if "\;q_1<0, q_2<0" solution will be similar)

Electric force: "\;\; F_2=k\frac{q_1 q_2 }{r^2}, \quad k=\frac{1}{4\pi \varepsilon _0}"

Magnetic force:"\;\;F_1=q_2 v_2 B_1\sin\theta, \quad \theta =90^{\circ}" angle between "v_2" and "B_1"

"\;\;F_1=q_2v B_1=q_2v\frac{\mu _0}{4\pi}\frac{q_1v}{r^2}=\frac{\mu _0}{4\pi}\frac{q_1 q_2 v^2}{r^2}"


"\frac{F_1}{F_2}=\mu_0 \varepsilon_0 v^2 =\lbrack \; c^2=\frac{1}{\varepsilon_0 \mu_0} \; \rbrack= \frac{v^2}{c^2}"

"\frac{F_1}{F_2}= \frac{(150\times 10^3)^2}{8.98\times 10^{16}}=25\times 10^{-8}"


Answer: "25\times 10^{-8}."